OpenStudy (anonymous):

what are the non-permissible values for the rational expression x(x+2)/(x-3)(x+1)

OpenStudy (anonymous):

\(\Huge\tt\color{red}{Welcome~}\color{orange}{To~}\color{yellow}{OpenStudy}!!\)

OpenStudy (xapproachesinfinity):

what non permissible mean actually? i wonder

OpenStudy (anonymous):

Whenever you use a rational expression, you must identify any values that must be excluded or considered non-permissible values. Non-permissible values are values that make the denominator

OpenStudy (xapproachesinfinity):

I'm googling since honestly never heard such terminology

OpenStudy (xapproachesinfinity):

oh i didn't see that you are talking about rational here i saw at as just a regular poly

OpenStudy (xapproachesinfinity):

excluded values you just solve for the bottom =0

OpenStudy (anonymous):

no they =1

OpenStudy (xapproachesinfinity):

no my dear values that are not allowed should bot equal to 0 that is why we set the bottom =0

OpenStudy (xapproachesinfinity):

(x+1)(x-3)=0 then x+1=0 or x-3=0

OpenStudy (anonymous):

no =1

OpenStudy (xapproachesinfinity):

so just solve for those to get the needed values

OpenStudy (anonymous):

no

OpenStudy (anonymous):

ghfghg

OpenStudy (anonymous):

hfgh

OpenStudy (anonymous):

fh

OpenStudy (anonymous):

h

OpenStudy (anonymous):

gfh

OpenStudy (anonymous):

g

OpenStudy (xapproachesinfinity):

don't listen to jojo ^_- the values that are exculded are x=-1 and x=3

OpenStudy (anonymous):

jojo please leave

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

okay that must be right because it is on of the multiple choice answers thanks

OpenStudy (anonymous):

d

OpenStudy (anonymous):

g

OpenStudy (anonymous):

g

OpenStudy (anonymous):

g

OpenStudy (anonymous):

g

OpenStudy (anonymous):

g

OpenStudy (anonymous):

gg

OpenStudy (anonymous):

g

OpenStudy (anonymous):

g

OpenStudy (xapproachesinfinity):

it must be right not because it is on your choices rather because we followed the principals that lead to good valid result ^_-

OpenStudy (anonymous):

i see

OpenStudy (xapproachesinfinity):

rational expressions were are not allowed to devide by 0 that's the meaning of excluding what gives 0 on bottom we don't have a meaning for something over 0

OpenStudy (xapproachesinfinity):

don't follow the answers, follow the logic and rules :)

OpenStudy (xapproachesinfinity):

that was the point i'm making

OpenStudy (anonymous):

@jojo100 please do not spam

OpenStudy (anonymous):

what i didnt and quit being so nice

OpenStudy (anonymous):

Who me?

OpenStudy (anonymous):

yea plz this plz that i just put a millon fn post get pissed

OpenStudy (anonymous):

Ive already been "rude" to 34 OS users, ive been suspended 22 times for it and warned 34 times so yea i dont want to get suspended again

OpenStudy (anonymous):

lol sorry for being a bick bro